OpenStudy (anonymous):

Taye and Mike are raking leaves to earn some money. Taye can rake 2 lawns that are about the same size in 3 hours. Mike can rake 2 lawns that are about the same in 4 hours. How long would it take both boys to work together to rake 2 lawns? Write and solve an equation for this situation. Explain how to set up the equation, using w = rt.

OpenStudy (anonymous):

idk it doesnt say

OpenStudy (paxpolaris):

ok .... \[\large w= r \times t\] work done= rate at which you work \(\times\) time you work for

OpenStudy (paxpolaris):

let Mike's Rate be M lawns per hour Taye's Rate be T lawns per hour

OpenStudy (paxpolaris):

\[\large 2= M \times 3\ \ \implies M=\frac 23\]\[\large 2= T \times 4\ \ \implies T=\frac 24=\frac 12\]

OpenStudy (paxpolaris):

when they both work together: w=rt \[\Large 2=\left( M+T\right) t\] where t is the time it takes when they work together

OpenStudy (paxpolaris):

\[\Large 2=\left( \frac 23+\frac 12\right) t\]

OpenStudy (anonymous):

thx man

OpenStudy (paxpolaris):

no problem.... \[t=1\frac57 hours\]